Grade 12Physical Chemistry

The ksp of caco3 and CaC2O4 as 7×10`-10 and 13,,× 10`-10washed with water.calculate the concentration of calcium ion

A Harish Kumar
9 Years agoGrade 12

1 Answer

Vikas TU
9 Years ago
       CaCO3       ===>           CaO      +  CO2
                                                s              s
ksp = s^2 = 7 x 10^-10
s = root(7) x 10^-5
 
For CaC2O4,
CaC2O4       =========>    Ca++(aq) + C2O4--(aq)  
                                              s                   s
 
Ksp = 13 x 10^-10 = s^2
s = root(13) x 10^-5
Total Concentration = [root(13) +  root(7)]*10^-5